OpenStudy (anonymous):

What is the surface area of a cylinder if the volume is 2 meters cubed and the height is 5 times the radius?

OpenStudy (campbell_st):

you need to 1st off find the radius... \[V = \pi r^2 h\] using h = 5r you end up with \[2 = \pi r^2 \times 5h .... so..... 2 = 5\pi r^3\] solve for r. then you will be able to find h which is 5r then find the surface area

hartnn (hartnn):

or since u have 2=5 pi r^3 keep (pi r^3)= 2/5 volume = 4/3 (pi r^3) = (4/3)*(2/5)=?